# Adversarial Reasoning for Repair Based on Inferred Program Intent

### v1

@@ Abstract.
AdverIntent-Agent was evaluated on two benchmarks: Defects4J 2.0 and HumanEval-Java. AdverIntent-Agent correctly repaired 77 and 105 bugs in both benchmarks, respectively. Our work suggests a new way for developers to decide whether to accept a patch by providing a complete package of program intent, tests, and code changes. This helps reduce the effort needed to review patches, going beyond just code transformation and test adequacy.
@@ 1. Introduction
Automated program repair (APR) aims to reduce the manual and costly processes involved in software maintenance tasks related to bug fixing (Monperrus2015, ; TSE-repair-survey, ). A wide range of approaches have been proposed to modify programs at the source level to bring behavior in line with a given test suite. These techniques include search-based (LeGoues2012GenProg, ; ssFix, ; sofix, ; Yuan2017ARJAAR, ; astor, ; ali-issta19-bytecode, ; sharpFix, ; relifix-shinwei-icse15, ; APR-Fuzzing-ISSTA23, ), semantics-based (Angelixicse16, ; acs, ; JFix, ; concolic-repair-PLDI21, ; patch-transplantation-TOSEM21, ; CrashProgramRepair-ISSTA19, ), and learning-based (codit-tse20, ; SEQUENCER, ; RewardRepair-icse22, ; Tufano-ICSE19, ; modit-ase21, ; CoCoNuT, ; Recoder, ; ye2022selfapr, ; ITER, ) techniques. Not unsurprisingly, the emergence of Large Language Models (LLMs) has motivated growing interest in using program repair techniques, such as code generated by GitHub Copilot (CopilotingCopilots, ), Amazon Code Whisperer (yeticstiren2023evaluating, ), and Replit (fan-automatedicse23, ). Recent large-scale industrial deployments from Google (googlerepair, ) and Meta (sapfix, ) provide promising initial validation that techniques developed in the lab can fruitfully move to industrial practice.
@@
However, existing deployments (and research prototypes) remain limited in the types of bugs they successfully target. Meta’s SapFix, for example, specifically targets certain types of null pointer exceptions. This allows tool designers to have more confidence that narrowly tailored patches are likely to be correct and acceptable. Acceptability in program repair has long been informed by program behavior on test suites, as tests are common in practice, and well-understood. However, overfitting to provided tests (cure-worse-15, ) is a known problem, resulting in patches that cause tests to pass, but do not generalize to the true specification, or developer intent. Researchers have proposed to tackle overfitting in traditional program repair through post-assessment approaches, such as additional heuristics (ICSE18-patchsim, ), probability models (ye2021ods, ; TianASE20, ; Jooyong-Poracle-TOMSE23, ), or test augmentation techniques (drr, ; gao2019crash, ). However, the problem is intrinsic in patch generation, and existing techniques remain limited in the complexity of defects they can empirically correctly tackle.
@@
We argue that such traditional techniques are limited in particular by their focus on transformation types in their design (such as new templates, or new types of synthesis, to use to attempt various types of repair). They do not, by and large, attempt to reason about higher level intent of a given program. An intent in this work is defined as the expected behavior of a function specification that developers aim to achieve. LLM-driven techniques have recently offered an alternative perspective, potentially increasing APR expressiveness and efficiency by both (a) engaging an LLM to reason about a program’s intended behavior, bringing developer intent more explicitly into the process, and (b) iteratively refining initial patch attempts towards more complex patches. This line of LLM-based work enables a self-repair loop to refine previously generated intermediate patches, which is considered state-of-the-art in the field (hidvégi2024cigar, ; ChatRepair-ISSTA24, ; ITER, ).
@@
This idea is promising, but incomplete: iterative, conversational repair approaches like ChatRepair (ChatRepair-ISSTA24, ), ITER (ITER, ), and Cigar (hidvégi2024cigar, ) fundamentally assume that an initial repair attempt proposed by an LLM can be refined to an acceptable solution. However, LLMs often face challenges in reasoning, and the initial bug-fixing intent is not always correct (llm-cannot-self-coorrect, ). In such cases, increasing the number of conversational iterations is unlikely to help (is-selfrepair-a-silver-bullet, ). Indeed, during such iterative processes, these approaches often cause conversational agents to issue refusal responses (wang-acl-2023-know, ; xie2024sorrybenchsystematicallyevaluatinglarge, ).
@@
Uses adversarial reasoning and testing to construct and identify the most likely correct intents. AdverIntent-Agent uses LLMs to infer multiple possible developer intentions with respect to a buggy function. These program intents are, by construction, intended to be independent and adversarial: that is, ideally, it should be impossible for the function, when repaired, to satisfy more than one. AdverIntent-Agent uses test generation to help demonstrate and validate that the inferred intentions are likely adversarial, assuming the intention is correct by construction and thereby sidestepping the oracle problem (oracleproblem, ). These tests allow for more effective automated reasoning about which intention is correct, and reduces the risk of overfitting to the developer-provided tests. The approach also increases the diversity of the considered patch pool by construction, increasing the likelihood of repair success. In practice, it is difficult to construct completely adversarial intents. Our work is an early exploration that uses test cases to quantify the degree of adversariality and to maximize the distinctiveness of intents within a limited number of attempts.
Dynamically constructs precise prompts to explore alternative bug root causes. AdverIntent-Agent uses adversarial reasoning to explore multiple possible program intentions. However, multiple root causes are still possible, even given a single possible program intent. AdverIntent-Agent uses dynamic precise prompts to first ask LLMs to reason about the top K root causes of bugs based on a given inferred intent. We use those root causes to seed additional prompts to request patches that address each root cause. This approach contrasts with previous uses of generic prompts for patch refinement (ChatRepair-ISSTA24, ), as it explicitly guides the generation of diverse solutions.
We implement AdverIntent-Agent and evaluate it against two benchmarks: Defects4J 2.0 (Defects4J, ) and HumanEval-Java (jiang2023knod, ). On the Defects4J 2.0 benchmark, AdverIntent-Agent successfully repairs the highest number of 77 bugs compared to related works (ChatRepair-ISSTA24, ; alpha-repair-fse22, ; repairagent, ; ye2022selfapr, ), and achieving a total of 105 bug repairs in HumanEval-Java, which surpasses all considered related works. These findings highlight the effectiveness of AdverIntent-Agent in addressing the challenges of automated program repair.
@@
  • Our work fundamentally shifts the interaction with developers regarding patch acceptance by providing a rich suggestion package that includes program intent, tests, and patches.
@@ 2. Illustration
((c)) Agentreason reasons three program intents
@@
((e)) Agentrepair generates three sets of patches to satisfy program intents and pass adversarial tests.
In this section, we illustrate the workflow of AdverIntent-Agent with an example. Consider 1(e), showing a bug in the CountUpper program from the Humaneval-Java (jiang2023knod, ) dataset. We use this example to demonstrate AdverIntent-Agent’s initial and adversarial approach to formulate program intention as part of a repair process.
Given the buggy program 1(e) and a test failure that exposes the bug 1(e), AdverIntent-Agent works as follows. First, a reasoning agent (Agentreason) starts with an initial prompt to reason about the intent of program, and its faulty statements: Summarize the program intent, and identify faulty locations. This results in the responses in 1(e). This produces an initial inferred intent and potentially buggy lines (first box of 1(e)), which pinpoints the for loop and the if condition are both incorrect, as it only checks every other letter with mixed lowercase and upper case vowels.
Agentreason also constructs an adversarial prompt. This allows AdverIntent-Agent to simultaneously explore a second possible intent. It therefore asks the LLM, “If the previous intent is NOT correct,…”, producing the adversarial intent to count uppercase vowels at every other position (show in the second box of 1(e)). This second prompt identifies the if condition on line 4 as problematic. AdverIntent-Agent continues this process to achieve another adversarial program intent to count both uppercase and lowercase vowels at even indices (box 3). Note that this differentiates our approach from prior agent-based work (repairagent, ; jiang2023knod, ) by considering different possibilities for program intents.
Next, 1(e) illustrates how the testing agent, Agenttest, uses the two adversarial intents to generate new differential test cases that reflect them. The goal of Agenttest is twofold. First, we use tests to measure the degree of adversarial behavior in the inferred program intents. The prompt instructs the LLM to construct tests with the same inputs but expected different outputs for each program intent. A greater number of differentially generated tests based on the same inputs indicates a higher degree of adversarial behavior in the program intents. When the number of tests falls below a predefined threshold, we reject and re-generate the inferred program intent to ensure all intents are adversarial to each other. In our example in 1(e), the adversarial score between the first and second intents is 50%, as two tests produce different results out of four total tests. In contrast, the adversarial score between intent 1 and intent 3 is 75%. Under our settings, all three intents are considered adversarial, as the threshold is set to 33.3% (the metric is introduced in subsection 3.2. Otherwise, the intents will be re-generated until they achieve the expected degree of adversarialness.
@@
The repair agent, Agentrepair, takes the output of both other agents to generate different sets of patches in 1(e). All generated patches are validated by both the original test cases to assess their plausibility, and on the newly generated adversarial tests reflecting inferred program intent. Our observation is that since the patches are adversarial, it is unlikely that all of them will pass the original test suite. Although incomplete, the developer-provided tests are therefore useful for testing adversarial patch plausibility. In our example, the second patch reflecting adversarial intent 2 is correct, while the other two are either overfitting or non-plausible. If only one is a plausible patch, this increases confidence in the inferred intent, and argues for the use of the new tests to supplement the test suite moving forward. If, on the other hand, all patches are plausible, AdverIntent-Agent increases the likelihood that at least one of them will be correct due to adversarial.
@@ 3. Approach
Figure 2. Overview of AdverIntent-Agent that shows three agents that are responsible for reasoning -Agentreason, test generation - Agenttest, and patch generation - Agentrepair.
@@
A core element of AdverIntent-Agent’s approach, as illustrated in 1(e), is the use of adversarial reasoning. The key to achieving reasoning is through criticism prompts to tell LLMs the previous answer is insufficient and explicitly to explore diversity, e.g., The previous answer is not correct, consider alternatives. Each agent thus simultaneously explores solutions to the initial repair task, as well as different adversarial formulations of the task. This approach is novel compared to related work as follows: 1) AdverIntent-Agent is the first work to reason about multiple, partially adversarial program intents, increasing the likelihood that at least one aligns with the correct intent. 2) AdverIntent-Agent is the first work, to our knowledge, to augment test generations during patch generation process, on the opposite of post patch assessments (ICSE18-patchsim, ; Jooyong-Poracle-TOMSE23, ) or specific oracle construction (e.g., crash (gao2019crash, ). 3) AdverIntent-Agent is explicitly guided to explore differences and alternatives, especially compared with prior conversational approaches (ChatRepair-ISSTA24, ; ITER, ), that limited themselves to iteratively refine generated patches on previous answers.
@@ 3.1. Agentreason
Table 1. Prompts used in AdverIntent-Agent
@@
Adversarial fault localization refined from the class level to statement granularity. Agentreason adversarially explores potential fault localizations by critiquing previous answers and guiding alternative exploration through three adversarial prompts. As shown in the first row of Table 1, Agentreason iteratively narrows the scope of faulty statements from the entire buggy class to finer granularity levels, such as functions (P1). Additionally, Agentreason explicitly guides LLMs to explore statements outside the function level, including constructors and variables, to achieve statement-level fault localization (P2). The LLM is also directed to examine related statements from the initially identified buggy class, as similar issues may exist in other classes, thereby enabling multi-location reasoning (P3). Initially, only a single class file is provided. After adversarial prompting and evaluating the error message, the LLM may request additional information. We then supply the requested files to facilitate cross-class fault localization. This design explicitly and systematically explores a diverse set of faulty statements, complementing previous answers. Note that Agentreason is novel in integrating LLM-based fault localization with program repair and reasoning about program behavior. Prior work either assumed perfect fault localization (e.g., RepairAgent (repairagent, )) and others  (ChatRepair-ISSTA24, ; alpha-repair-fse22, ; CURE-icse21, ; SEQUENCER, ; CoCoNuT, ). Other techniques integrate with spectrum-based fault localization (see, e.g., GenProg (le_goues_genprog_2012, ) and related techniques  (RewardRepair-icse22, ; ye2022selfapr, )). This includes agent-like AutoCodeRover (autocoderover, ), which considers fault localization and repair as two separate tasks and increases configuration effort.
@@
With the identified faulty functions and statements from the first step, Agentreason begins by generating the initial program intent by asking the LLM, ”Can you reason about the program intent and locate the corresponding fault statements?” (P4). Based on the response, Agentreason explicitly guides the LLM to reason about alternative program intents by posing the critical prompt, ”What if the previous intents are incorrect? What alternatives are available?” (P5). These critical prompts can be repeated K times until K adversarial intents are generated. An example is provided earlier in 1(e), where the response summarizes the program intents in natural language along with the corresponding faulty statements that violate those intents. Our insights are twofold. First, Agentreason provides multiple program intents, which are, to some degree, adversarial to each other, thus covering the entire search space of possible intents. This increases the likelihood that at least one of them is correct or close to the original developer’s intent. Second, by assuming each individual intent is correct, this approach partially addresses the oracle problem, enabling test oracles to be constructed as long as they satisfy the inferred program intent.

### v2

@@ Abstract.
AdverIntent-Agent was evaluated on two benchmarks: Defects4J 2.0 and HumanEval-Java. AdverIntent-Agent correctly repaired 77 and 105 bugs in both benchmarks, respectively. Our work helps reduce the effort required to review patches by enabling developers to assess program intent in natural language, rather than reviewing code patches.
@@ 1. Introduction
Automated program repair (APR) aims to reduce the manual and costly processes involved in software maintenance tasks related to bug fixing (Monperrus, 2017; Gazzola et al., 2017). A wide range of approaches have been proposed to modify programs at the source level to bring behavior in line with a given test suite. These techniques include search-based (Le Goues et al., 2012a; Xin and Reiss, 2017b; Liu and Zhong, 2018; Yuan and Banzhaf, 2018; Martinez and Monperrus, 2016a; Ghanbari et al., 2019; Xin and Reiss, 2019; Tan and Roychoudhury, 2015; Kim et al., 2023), semantics-based (Mechtaev et al., 2016; Xiong et al., 2017; Le et al., 2017; Shariffdeen et al., 2021a, b; Gao et al., 2019a), and learning-based (Chakraborty et al., 2020; Chen et al., 2019; Ye et al., 2022a; Tufano et al., 2019; Chakraborty and Ray, 2021; Lutellier et al., 2020; Zhu et al., 2021; Ye et al., 2022b; Ye and Monperrus, 2024) techniques. Not unsurprisingly, the emergence of Large Language Models (LLMs) has motivated growing interest in using program repair techniques, such as code generated by GitHub Copilot (Wei et al., 2023), Amazon Code Whisperer (Yetiştiren et al., 2023), and Replit (Fan et al., 2023). Recent large-scale industrial deployments from Google (Google, 2024) and Meta (Marginean et al., 2019) provide promising initial validation that techniques developed in the lab can fruitfully move to industrial practice.
@@
However, existing deployments (and research prototypes) remain limited in the types of bugs they successfully target. Meta’s SapFix, for example, specifically targets certain types of null pointer exceptions. This allows tool designers to have more confidence that narrowly tailored patches are likely to be correct and acceptable. Acceptability in program repair has long been informed by program behavior on test suites, as tests are common in practice, and well-understood. However, overfitting to provided tests (Smith et al., 2015) is a known problem, resulting in patches that cause tests to pass, but do not generalize to the true specification, or developer intent. Researchers have proposed to tackle overfitting in traditional program repair through post-assessment approaches, such as additional heuristics (Xiong et al., 2018), probability models (Ye et al., 2021a; Tian et al., 2020; Ismayilzada et al., 2023), or test augmentation techniques (Ye et al., 2021b; Gao et al., 2019b). However, the problem is intrinsic in patch generation, and existing techniques remain limited in the complexity of defects they can empirically correctly tackle.
@@
We argue that such traditional techniques are limited in particular by their focus on transformation types in their design (such as new templates, or new types of synthesis, to use to attempt various types of repair). They do not, by and large, attempt to reason about higher level intent of a given program. An intent in this work is defined as the expected behavior of a function specification that developers aim to achieve. LLM-driven techniques have recently offered an alternative perspective, potentially increasing APR expressiveness and efficiency by both (a) engaging an LLM to reason about a program’s intended behavior, bringing developer intent more explicitly into the process, and (b) iteratively refining initial patch attempts towards more complex patches. This line of LLM-based work enables a self-repair loop to refine previously generated intermediate patches, which is considered state-of-the-art in the field (Hidvégi et al., 2024; Xia and Zhang, 2024; Ye and Monperrus, 2024).
@@
This idea is promising, but incomplete: iterative, conversational repair approaches like ChatRepair (Xia and Zhang, 2024), ITER (Ye and Monperrus, 2024), and Cigar (Hidvégi et al., 2024) fundamentally assume that an initial repair attempt proposed by an LLM can be refined to an acceptable solution. However, LLMs often face challenges in reasoning, and the initial bug-fixing intent is not always correct (Huang et al., 2024). In such cases, increasing the number of conversational iterations is unlikely to help (Olausson et al., 2024). Indeed, during such iterative processes, these approaches often cause conversational agents to issue refusal responses (Wang et al., 2023; Xie et al., 2024).
@@
Uses adversarial reasoning and testing to construct and identify the most likely correct intents. AdverIntent-Agent uses LLMs to infer multiple possible developer intentions with respect to a buggy function. These program intents are, by construction, intended to be independent and adversarial: that is, ideally, it should be impossible for the function, when repaired, to satisfy more than one. AdverIntent-Agent uses test generation to help demonstrate and validate that the inferred intentions are likely adversarial, assuming the intention is correct by construction and thereby sidestepping the oracle problem (Barr et al., 2015). These tests allow for more effective automated reasoning about which intention is correct, and reduces the risk of overfitting to the developer-provided tests. The approach also increases the diversity of the considered patch pool by construction, increasing the likelihood of repair success. In practice, it is difficult to construct completely adversarial intents. Our work is an early exploration that uses test cases to quantify the degree of adversariality and to maximize the distinctiveness of intents within a limited number of attempts.
AdverIntent-Agent uses adversarial reasoning to explore multiple possible program intentions. However, multiple root causes are still possible, even given a single possible program intent. AdverIntent-Agent uses dynamic precise prompts to first ask LLMs to reason about the top-k root causes of bugs based on a given inferred intent. We use those root causes to seed additional prompts to request patches that address each root cause. This approach contrasts with previous uses of generic prompts for patch refinement (Xia and Zhang, 2024), as it explicitly guides the generation of diverse solutions.
We implement AdverIntent-Agent and evaluate it against two benchmarks: Defects4J 2.0 (Just et al., 2014) and HumanEval-Java (Jiang et al., 2023b). On the Defects4J 2.0 benchmark, AdverIntent-Agent successfully repairs 77 bugs, outperforming related works (Xia and Zhang, 2024, 2022; Bouzenia et al., 2024; Ye et al., 2022b). In HumanEval-Java, AdverIntent-Agent achieves 105 bug repairs, also surpassing the considered related works. These findings highlight the effectiveness of AdverIntent-Agent in addressing the challenges of automated program repair.
@@ 2. Illustration
((c)) Agentreason reasons three program intents.
@@
((e)) Agentrepair generates three sets of patches to satisfy program intents and pass adversarial tests
In this section, we illustrate the workflow of AdverIntent-Agent with an example. Consider Figure 1, showing a bug in the CountUpper program from the Humaneval-Java (Jiang et al., 2023b) dataset. We use this example to demonstrate AdverIntent-Agent’s initial and adversarial approach to formulate program intention as part of a repair process.
Given the buggy program and a test failure that exposes the bug, as shown in Figure 1, AdverIntent-Agent works as follows. First, a reasoning agent (Agentreason) starts with an initial prompt to reason about the intent of program, and its faulty statements: Summarize the program intent, and identify faulty locations. These responses are shown in Figure 1. This produces an initial inferred intent and potentially buggy lines (first box of Figure 1), which pinpoints the for loop and the if condition are both incorrect, as it only checks every other letter with mixed lowercase and upper case vowels.
Agentreason also constructs an adversarial prompt. This allows AdverIntent-Agent to simultaneously explore a second possible intent. It therefore asks the LLM, “If the previous intent is NOT correct,…”, producing the adversarial intent to count uppercase vowels at every other position (shown in the second box of Figure 1). This second prompt identifies only the if condition on line 4 as problematic. AdverIntent-Agent continues this process to achieve another adversarial program intent to count both uppercase and lowercase vowels at even indices (shown in the third box of Figure 1). Note that this differentiates our approach from prior agent-based work (Bouzenia et al., 2024; Jiang et al., 2023b) by considering different possibilities for program intents.
Next, Figure 1 illustrates how the testing agent, Agenttest, uses the two adversarial intents to generate new differential test cases that reflect them. The goal of Agenttest is twofold. First, we use tests to measure the degree of adversarial behavior in the inferred program intents. The prompt instructs the LLM to construct tests with the same inputs but expected different outputs for each program intent. A greater number of differentially generated tests based on the same inputs indicates a higher degree of adversarial behavior in the program intents. When the number of tests falls below a predefined threshold, we reject and re-generate the inferred program intent to ensure all intents are adversarial to each other. In our example in Figure 1, the adversarial score between the first and second intents is 50%, as two tests produce different results out of four total tests. In contrast, the adversarial score between intent 1 and intent 3 is 75%. Under our settings, all three intents are considered adversarial, as the threshold is set to 33.3% (the metric is introduced in subsection 3.2. Otherwise, the intents will be re-generated until they achieve the expected degree of adversarialness.
@@
The repair agent, Agentrepair, takes the output of both other agents to generate different sets of patches in Figure 1. All generated patches are validated by both the original test cases to assess their plausibility, and on the newly generated adversarial tests reflecting inferred program intent. Our observation is that since the patches are adversarial, it is unlikely that all of them will pass the original test suite. Although incomplete, the developer-provided tests are therefore useful for testing adversarial patch plausibility. In our example, the second patch reflecting adversarial intent (middle box of Figure 1) is correct, while the other two are either overfitting (first box of Figure 1) or non-plausible (third box of Figure 1). If only one is a plausible patch, this increases confidence in the inferred intent, and argues for the use of the new tests to supplement the test suite moving forward. If, on the other hand, all patches are plausible, AdverIntent-Agent increases the likelihood that at least one of them will be correct due to adversarial.
@@ 3. Approach
Figure 2. Overview of AdverIntent-Agent that shows three agents that are responsible for reasoning -Agentreason, test generation - Agenttest, and patch generation - Agentrepair. AdverIntent-Agent begins with a buggy class and failing tests as inputs, and its outputs include new intents, tests, and candidate patches.
@@
A core element of AdverIntent-Agent’s approach, as illustrated in the motivating example, is the use of adversarial reasoning. The key to achieving reasoning is through criticism prompts to tell LLMs the previous answer is insufficient and explicitly to explore diversity, e.g., The previous answer is not correct, consider alternatives. Each agent thus simultaneously explores solutions to the initial repair task, as well as different adversarial formulations of the task. This approach is novel compared to related work as follows: 1) AdverIntent-Agent is the first work to reason about multiple, partially adversarial program intents, increasing the likelihood that at least one aligns with the correct intent. 2) AdverIntent-Agent is the first work, to our knowledge, to augment test generations during patch generation process, on the opposite of post patch assessments (Xiong et al., 2018; Ismayilzada et al., 2023) or specific oracle construction (e.g., crash (Gao et al., 2019b). 3) AdverIntent-Agent is explicitly guided to explore differences and alternatives, especially compared with prior conversational approaches (Xia and Zhang, 2024; Ye and Monperrus, 2024), that limited themselves to iteratively refine generated patches on previous answers.
@@ 3.1. Agentreason
Table 1. Prompts used in AdverIntent-Agent.
@@
Adversarial fault localization refined from the class level to statement granularity. Agentreason adversarially explores potential fault localizations by critiquing previous answers and guiding alternative exploration through three adversarial prompts. As shown in the first row of Table 1, Agentreason iteratively narrows the scope of faulty statements from the entire buggy class to finer granularity levels, such as functions (P1). Additionally, Agentreason explicitly guides LLMs to explore statements outside the function level, including constructors and variables, to achieve statement-level fault localization (P2). The LLM is also directed to examine related statements from the initially identified buggy class, as similar issues may exist in other classes, thereby enabling multi-location reasoning (P3). Initially, only a single class file is provided. After adversarial prompting and evaluating the error message, the LLM may request additional information. We then supply the requested files to facilitate cross-class fault localization. This design explicitly and systematically explores a diverse set of faulty statements, complementing previous answers. Note that Agentreason is novel in integrating LLM-based fault localization with program repair and reasoning about program behavior. Prior work either assumed perfect fault localization (e.g., RepairAgent (Bouzenia et al., 2024)) and others  (Xia and Zhang, 2024, 2022; Jiang et al., 2021; Chen et al., 2019; Lutellier et al., 2020). Other techniques integrate with spectrum-based fault localization (see, e.g., GenProg (Le Goues et al., 2012b) and related techniques  (Ye et al., 2022a, b)). This includes agent-like AutoCodeRover (Zhang et al., 2024b), which considers fault localization and repair as two separate tasks and increases configuration effort.
@@
With the identified faulty functions and statements from the first step, Agentreason begins by generating the initial program intent by asking the LLM, ”Can you reason about the program intent and locate the corresponding fault statements?” (P4). Based on the response, Agentreason explicitly guides the LLM to reason about alternative program intents by posing the critical prompt, ”What if the previous intents are incorrect? What alternatives are available?” (P5). These critical prompts can be repeated K times until K adversarial intents are generated. An example is provided earlier in Figure 1, where the response summarizes the program intents in natural language along with the corresponding faulty statements that violate those intents. Our insights are twofold. First, Agentreason provides multiple program intents, which are, to some degree, adversarial to each other, thus covering the entire search space of possible intents. This increases the likelihood that at least one of them is correct or close to the original developer’s intent. Second, by assuming each individual intent is correct, this approach partially addresses the oracle problem, enabling test oracles to be constructed as long as they satisfy the inferred program intent.